OpenStudy (anonymous):

PLEASE HELP ALGEBRA 2 QUESTION (EASY!) The perimeter of a rectangular book cover is 106 cm. The width is 11 cm more than half the length? What is the length of the book cover?

OpenStudy (yb1996):

2x + 2(x+11) = 106 2x + 2x + 22 = 106 4x = 84 x = 21 The length is 21 cm.

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

But wait that makes no sense...

OpenStudy (anonymous):

or wait sorry

OpenStudy (anonymous):

yes it makes 83, not 106

OpenStudy (yb1996):

Oh gosh, sorry, I read the problem wrong. I'll fix it for you, give me s sec.

OpenStudy (anonymous):

it's fine :D thank you!

OpenStudy (yb1996):

2x + 2( 1/2 + 11) = 106 2x + x + 22 = 106 3x = 84 x = 28

OpenStudy (anonymous):

let me check xD

OpenStudy (anonymous):

thank you so much for your help!

OpenStudy (yb1996):

no problem ..... sorry, the first line is supposed to be "2x + 2( 1/2 x + 11) = 106" I forgot to put an x there

OpenStudy (anonymous):

Thank you so much! It's right :)